OpenStudy (amonoconnor):

When dealing with Deriviatives, and pulling information of an function's graph from said function analytically, if the 2nd derivative set equal to 0 is undefined/irrational, does that mean there is an asymptote along the graph? Specifically, the problem I'm working on has a second derivative of "2 + 6x^(-4)". So, theoretically, x = (-3)^(1/4). This is undefined/ irrational, yet what does this mean? How do we know which values of x have the asymptote, IF that's what an undefined answer here means? :/ I need help with understanding on this! Any and all help is greatly appreciated! :)

OpenStudy (amonoconnor):

Also, does this mean that the graph of the function has no Inflection Points?

OpenStudy (anonymous):

what is the original function?

OpenStudy (amonoconnor):

y = ((x^4) + 1)/x^2

OpenStudy (anonymous):

better known as \(f(x)=x^2+\frac{1}{x^2}\) ok

OpenStudy (amonoconnor):

Yep, I get that. :)

OpenStudy (anonymous):

so the second derivative is \[\frac{6}{x^4}+2\]right?

OpenStudy (amonoconnor):

Yes, that's what I got.

OpenStudy (anonymous):

or as you had it \[2+6x^{-4}\]

OpenStudy (amonoconnor):

:)

OpenStudy (anonymous):

the problem with using exponential notation with negative exponents is that you cannot really compute with them or solve with them

OpenStudy (anonymous):

perhaps it was not clear how to solve \[2+6x^{-4}=0\] but it is more clear with \[2+\frac{6}{x^2}=0\] or at least clearer

OpenStudy (anonymous):

or rather \[2+\frac{6}{x^4}=0\] it is not obvious that there is no solution

OpenStudy (anonymous):

i meant to say "it is NOW obvious that there is no solution"

OpenStudy (amonoconnor):

So... 2 + (6/x^4) = 0 -2 _________________ 6/(x^4) = -2 (x^4)(6/(x^4)) = -2(x^4) 6 = -2x^4 __________ -2 x^4 = -3 Yeah, get how that is found! but what, conceptually, does it MEAN in terms of the graph ??

OpenStudy (anonymous):

hold the phone

OpenStudy (amonoconnor):

:) sorry it took me a while to type out that math.

OpenStudy (amonoconnor):

Okay

OpenStudy (anonymous):

you are doing tons too much work whatever \(x\) is it is certain that \(x^4\) is positive, and since \(6\) is positive we know \(\frac{6}{x^4}\) is positive, so certainly \(2+\frac{6}{x^4}\) is positive and therefore it is not zero

OpenStudy (anonymous):

so there is no solution i noticed you wrote "irrational/undefined" it is not "irrational" there is no real number solution what this tells you is that the second derivative is always positive what that says about the original function is that is always leans left, or as your math teacher probably says "concave up"

OpenStudy (amonoconnor):

So... If the second derivative is not solvable for x (undefined) then the graph is always concave up, all the time?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

if it is never zero, then it is always either positive or negative (assuming it is continuous) that means the function is either always concave up or concave down, depending on whether the second derivative is positive or negative

OpenStudy (amonoconnor):

*rgggg* ;)

OpenStudy (amonoconnor):

Ahhhhh.... Okay. That makes beautiful, perfect sense.

OpenStudy (anonymous):

clear right? never zero doesn't mean always positive, just in this case it was it could also be always negative

OpenStudy (amonoconnor):

:))) thank you so much!

OpenStudy (anonymous):

think of a real simple example, like \(y=ax^2+bx+c\) where \(y''=2a\)

OpenStudy (anonymous):

if \(a>0\) then it is always concave up, which you knew already since it is a parabola that opens up but if \(a<0\) then it is always concave down

OpenStudy (anonymous):

yw

OpenStudy (amonoconnor):

a = 0 (?)

OpenStudy (amonoconnor):

So, the I.P. Is at x = 0, and it's concave up?

OpenStudy (anonymous):

then so is the second derivative, and you have a line

OpenStudy (anonymous):

if \(a=0\) then \(y=ax^2+bx+c=bx+c\) a line which is neither concave up or down

OpenStudy (amonoconnor):

Don't you plug the value of x from the 2nd deriviative into the original equation to find the y-value of the I.P.?

OpenStudy (anonymous):

i don't know what an IP is

OpenStudy (anonymous):

oh inflection point nvm

OpenStudy (amonoconnor):

Sorry

OpenStudy (anonymous):

if the second derivative is never zero, then there is not inflection point

OpenStudy (amonoconnor):

Okay

OpenStudy (anonymous):

no inflection point is what i was trying to write the inflection point is at the zero of the second derivative

OpenStudy (amonoconnor):

Yep, I think I get it all now. Ha, I know I at least get it a heck of a lot better. :)

OpenStudy (amonoconnor):

Is there an asymptote at x = 0 then?

OpenStudy (amonoconnor):

@satellite73

OpenStudy (anonymous):

for the original function?

OpenStudy (amonoconnor):

Yeah in the graph of the original function.

OpenStudy (anonymous):

yes, because that would make the denominator equal to zero

OpenStudy (amonoconnor):

Okay.

OpenStudy (amonoconnor):

So... are +/- 1 minima (1,2) & (-1.2) ?

OpenStudy (anonymous):

i don't know, i didn't do it hold on

OpenStudy (anonymous):

yes

OpenStudy (amonoconnor):

Alright.... And the , because it will approach 0, to infinity, it is increasing over the intervals: [-1, 0) U [1, 🔃)

OpenStudy (anonymous):

increasing on \((-1,0)\) and \((1\infty)\) looks good make the intervals open

OpenStudy (amonoconnor):

Okay, sweet. Awesome, this is great, I've learned more in like, an hour online than I have in two weeks in Calc at school.... That's nice. :p thank you :)

OpenStudy (anonymous):

your quite welcome good luck

OpenStudy (amonoconnor):

I have another problem, quite similar- would it be alright if I sort of continue into that one, and ask you questions again as they arise in my brain?

OpenStudy (amonoconnor):

Haha... I am sorry, and feel free to say no, but what would you say the second deriviative of "y = x/((x^2) - 4)" is?

OpenStudy (amonoconnor):

I used the quotient rule right out of the gate, is that what you'd do?

OpenStudy (anonymous):

\[\frac{x}{x^2-4}\]?

OpenStudy (amonoconnor):

Yep :)

OpenStudy (anonymous):

\[\frac{2x^3+24x}{(x-4)^3}\] i think

OpenStudy (anonymous):

typo there try \[\frac{2x^3+24x}{(x^2-4)^3}\]